OpenStudy (anonymous):

Solve the inequality. |r + 3| ≥ 7

OpenStudy (anonymous):

show work please.

OpenStudy (anonymous):

See, first equation becomes: \[r + 3 \ge 7\] Subtract -3 both the sides..

OpenStudy (anonymous):

\[r+3\ge 7 \] and \[r+3\le -7\]

OpenStudy (anonymous):

thus ,\[r \ge 4\] and \[r \le -10\]